write the equation of the line (in slope intercept form) that passes through the point (7,5) and is perpendicular to the line x-2y=7.
First I got y on the left side: y=x/2-7/2
Now I really don't know what to do. Please help.

Answer by stanbon(75887)

write the equation of the line (in slope intercept form) that passes through the point (7,5) and is perpendicular to the line x-2y=7.
First I got y on the left side: y=x/2-7/2
-----
Since y = (1/2)x -(7/2)
the slope of any line that is perpendicular to it
is -2.
---
Then you can find the equation you are looking for.
Form is y = mx + b
m = -2 and y = 5 when x = 7.
Substitute and solve for "b":
5 = -2(7) + b
5 = -14 + b
b = 19
----------
Equation is: y = -2x+19
==============================
Cheers,
Stan H.
==============================